OpenStudy (anonymous):

If the hypotenuse of a 45-45-90 triangle is 5 sqrt 3 how do I find the two legs?

OpenStudy (dmb1215):

oke how long is the hypotenuse

OpenStudy (dmb1215):

All i need is the hypotenuse then i will have the rest for you

OpenStudy (anonymous):

the hypotenuse is 5 square root three

OpenStudy (dmb1215):

ok

OpenStudy (dmb1215):

125 2√ 2 heres one

OpenStudy (dmb1215):

its the same for both

OpenStudy (anonymous):

I understand its the same but how do I get those two leg answers?

OpenStudy (anonymous):

as two sides have same angle (45,45) so these have the same length so let hypotenuse = z 1st leg = x 2nd leg = y from pythagoras theorem \[z^2 = x^2 + y^2\] as x = y {same angles} so \[z^2 = x^2 + x^2\] \[z^2 = 2x^2\] putting the values \[( 5 \sqrt{3} )^2 = 2x^2\] 25(3) = 2x^2 75 = 2x^2 75/2 = x^2 \[\sqrt{37.5} = x \] So \[1st leg = \sqrt{37.5} \] \[2nd leg = \sqrt{37.5} \]

OpenStudy (anonymous):

Wow. That's kinda confusing but I think I understand it a little bit better. Thank you very much